Find the distance if you travel at a rate of 65mph for 7 hours. Hint: Distance = Rate × Time
Answer by TimothyLamb(4379)

---
s = d/t
d = s*t
d = 65 * 7
d = 455 miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php